OpenStudy (anonymous):

Part 1: Solve each of the quadratic equations below. Show your work. (3 points) x2 – 36 = 0 and x2 = 8x – 12 Part 2: Describe what the solution(s) represent to the graph of each. (2 points) Part 3: How are the graphs alike? How are they different? (2 points)

OpenStudy (uri):

So do you have to solve 2 question? I'm talking about Part 1

OpenStudy (anonymous):

x^2 – 36 = 0 First, replace the "b" variable by 0 since there isn't any. Then find the factors of -36 that adds up to 0. x^2 + 0 - 36 = 0 1 * -36 2 * -18 3 * -12 4 * -9 6 * 6 -1 * 36 -2 * 18 -3 * 12 -4 * 9 -6 * 6 The one that adds up to 0 is 6 and -6 so now you factor normally by making them equal 0's. (x + 6)(x - 6) x+6 = 0 x-6 = 0 x= -6 and x= 6

OpenStudy (anonymous):

I was actually answering that question on an online test a few moment ago what a coincidence

OpenStudy (anonymous):

Unlike the others this one has x^2 in the beginning so you just subtract it from both sides. That way it'll equal 0 like a normal equation. x^2 = 8x – 12 0 = -x^2 + 8x - 12 Because the -x^2 is negative we can't go on, but if we multiply both sides by -1 and now we have all positive numbers, then we can go on and find factors of 12 that add up to 8. 0 = x^2 + 8x + 12 1 * 12 2 * 6 3 * 4 2 and 6 will do, now we factor normally and make them equal 0. (x + 2)(x + 6) x+2 = 0 x+6 = 0 x= -2 and x= -6

OpenStudy (anonymous):

Do you get the steps?

OpenStudy (anonymous):

@Rexy705 yeah I do.

OpenStudy (anonymous):

what the solution x's equal tell u were the parabola will intersect the x-axis